I need help please

Suppose the market demand and supply functions are QD = 2450 – 10P and QS = 15P + 100. You have just graduated and moved to this city; as a new MBA and an entrepreneur, you are considering entering the market for this product.

a.   Determine the equilibrium price and quantity in this market.

b.   You’ve researched and found that most firms in the market currently experience costs such that TC = 50 + 115Q – 4Q2 + 0.15Q3. Determine whether or not you should enter this market. Use graphs to support your answer. (Remember that you can Format Axis and change the Minimum and Maximum Bounds of your axes to “zoom in” to a graph in Excel.)

c.   Due to unforeseen delays, you don’t enter the market. However, a year later the market supply has changed to QS = 16P + 110. Are you surprised at this shift in supply?

d.   Given the new supply conditions, determine whether or not you should enter the market.
Answer #1

a.

The demand function is QD: 2450 – 10P

The supply function is: Qs=15P+100

At equilibrium, QD=QS

This implies,

QD=QS

2450-10P=15P+100

25P=2350

P=94

Hence, equilibrium price is P=94

Hence, equilibrium quantity is: Q=2450-10*94 = 2450-940=1510

Hence, equilibrium price is P=94

Equilibrium quantity is Q=1510

b.

The total cost is TC = 50+115Q-4Q^2+0.15Q^3

Hence, the fixed cost TFC = 50

The total variable cost is: TVC = 115Q-4Q^2+0.15Q^3

The marginal cost is: MC = 115-8Q+0.45Q^2

The given total cost function suggest that the industry is decreasing cost industry. Thus, the firm should enter the market.

Market Equilibrium Q AC AVC MC 94 1 161.15 111.15 107.2025 180 94 2 132.6 107.6 99.2025 94 3 121,0167 104.35 91.2025 160 4 94Thus, as can be observed, the marginal cost decreases with in creasing level of output produced. Hence, firm should enter the market.

c. The market supply changed to QS=16P+110

One should not be surprised. This is because, the over time there will be more and more firms that would enter the market. This is market by the rightward shift in the supply curve as indicated by shift in the constant from 100 to 110. With increasing supply, the competition would increase and price per unit of product would fall. This is market by the increase in the value of the coefficient of P in QS equation from 15 to 16

d.

Thus, the new market equilibrium is:

QD=QS

2450-10P=16P+110

26P=2340

P=90

At P=90, equilibrium quantity is: Q=2450-10*90 = 2450-900=1550